A vertical aerial AB 12.65m high stands on ground which is inclined 18°32 15to the horizontal. A stay connects the top of the aerial A to a point Con the ground 20.0 m uphill from Bthe foot of the aerial.

Determine:

the length of the stay .

the angle the stay makes with the ground


1
Expert's answer
2021-12-09T04:46:03-0500


Given "AB=12.65m, BC=20.0m, \alpha=18\degree32'15''."


"\beta=90\degree-\alpha"

Then


"\angle ABC=\beta=90\degree-\alpha"

Triangle ACB


"AB=12.65, BC=20, ''"

"\angle ABC=90\degree-\alpha=90\degree-18\degree32'15''"

a) Law of Cosines


"{AC}^2={AB}^2+{BC}^2-2(AB)(BC)\cos\angle ABC"

"AC=\sqrt{{12.65}^2+{20}^2-2(12.65)(20)\sin(18\degree32'15'')}"

"AC=19.98m"

2) Law of Sines


"\dfrac{\sin\angle ACB}{AB}=\dfrac{\sin \angle ABC}{AC}"

"\angle ACB=\sin^{-1}(\dfrac{AB\sin \angle ABC}{AC})"

"\angle ACB=\sin^{-1}(\dfrac{12.65\cos (18\degree32'15'')}{19.98})"

"\angle ACB=36.89\degree"

"\angle ACB=36\degree53'25''"